# Discovery and Expansion of New Domains within Diffusion Models

### v2

@@ Abstract
In this work, we study the generalization properties of diffusion models in a few-shot setup, introduce a novel tuning-free paradigm to synthesize the target out-of-domain (OOD) data, and demonstrate its advantages compared to existing methods in data-sparse scenarios with large domain gaps. Specifically, given a pre-trained model and a small set of images that are OOD relative to the model’s training distribution, we explore whether the frozen model is able to generalize to this new domain. We begin by revealing that Denoising Diffusion Probabilistic Models (DDPMs) trained on single-domain images are already equipped with sufficient representation abilities to reconstruct arbitrary images from the inverted latent encoding following bi-directional deterministic diffusion and denoising trajectories. We then demonstrate through both theoretical and empirical perspectives that the OOD images establish Gaussian priors in latent spaces of the given model, and the inverted latent modes are separable from their initial training domain. We then introduce our novel tuning-free paradigm to synthesize new images of the target unseen domain by discovering qualified OOD latent encodings in the inverted noisy spaces. This is fundamentally different from the current paradigm that seeks to modify the denoising trajectory to achieve the same goal by tuning the model parameters. Extensive cross-model and domain experiments show that our proposed method can expand the latent space and generate unseen images via frozen DDPMs without impairing the quality of generation of their original domain. We also showcase a practical application of our proposed heuristic approach in dramatically different domains using astrophysical data, revealing the great potential of such a generalization paradigm in data spare fields such as scientific explorations.
@@ 1 Introduction
Generalization ability, which enables the model to synthesize data from various domains, has long been a challenge in deep generative models. The current research trend focuses on leveraging larger models with more training data to facilitate improved generalization. The popularity of recent large-scale models such as DALLE-2 [40], Imagen [20] and StableDiffusion [43] have demonstrated the impressive and promising representation abilities of the state-of-the-art (SOTA) diffusion generative models when trained on an enormous amount of images such as LAION-5B [44]. However, scaling up is not a panacea and does not fundamentally solve the generalization challenge. In other words, for data domains that remain sparse in those already giant datasets, such as the astrophysical observation and simulation data, even the SOTA models fail to synthesize data suitable for rigorous scientific research purposes, as illustrated in the middle columns of Fig. 1. In addition, scaling up requires exhaustive resources, severely limiting the number of research groups that are able to participate and contribute to the work, and consequently hindering research progress. Given the aforementioned concerns, our work focuses on studying the generalization ability in a few-shot setup, where a pre-trained diffusion generative model and a small set of raw images different from its training domain are provided, with the ultimate objective of generating new data samples from the target OOD domain.
Figure 1: Examples of synthesized OOD images across data domains and methods. (Left): Using a pre-trained DDPM on dog images as an example, we can well reconstruct arbitrary unseen images across domains covering natural images and astrophysical data. (Middle): Even though large SOTA generative models have strong generalization abilities, their synthesized data are still very different from real data in specific domains. (Right): In this work, we propose a discovery-based tuning-free paradigm for synthesizing OOD data in a few-shot scenario. Different from tuning-based methods that usually yield better performance with closer domain gaps (e.g., dogs to humans), our method shows superiority as the domain gaps become larger.
Recall the generation process with a diffusion model (or most generative models in general), which essentially involves two steps: latent encoding sampling [^1] The latent spaces in this work are different from the concept of LDMs (aka, StableDiffusion [43]), where the diffusion denoising operates on the latent space of raw data from autoencoders. Our latent spaces refer to intermediate noisy states along the denoising chain, similar to [74]. and denoising. [^2] The denoising can be considered as the trajectory mapping between the latent space and the real data space. In the realm of natural images within the computer vision community, most existing works address the problem by modifying trajectories and fine-tuning model parameters using new images and extra semantic supervision from the text [29, 67, 31]. However, tuning-based methods have several drawbacks, especially given the large transition gap between trained and target data domains. It is known that tuning methods do not generalize well when the domain gap between the target OOD domain and the training domain becomes too large [69, 70, 56]. In fact, our experiments show that tuning pre-trained DMs using only image supervisions with vanilla variational lower bound loss [19] is very difficult. Only with additional semantic guidance, such as CLIP loss [39], can tuning performance improve, as demonstrated in Fig. 1. Moreover, modifying parameters adversely affects synthesis quality in the model’s originally trained domain. The tuning cost is also entirely dependent on the pre-trained model, which can be quite high given the well-known expensive training for diffusion models. Unlike the existing trajectory-tuning paradigm, we introduce a heuristic approach featuring a novel tuning-free paradigm that achieves the objective by sampling latent encodings of the unseen target domain within the latent spaces of pre-trained diffusion models. Our core idea is to discover the corresponding OOD latent encodings and then denoise them through deterministic trajectories, as depicted in Fig. 2.
We begin with an in-depth study of the generalization properties of diffusion models. As a pivotal observation, we note that the SOTA diffusion generative models pre-trained on single-domain images (e.g., dog faces) already have sufficient representation ability to reconstruct images from arbitrary unseen domains (e.g., human faces and churches) from the inverted OOD latent encodings, [^3] Inversion refers to the process of converting raw data to noisy encodings in the literature of generative modeling [61], which can also be understood as a diffusion process in the context of DMs. as qualitatively shown in Fig. 1, and with details explained in Sec. 2.2. The results from the arbitrary image reconstruction test suggest that an inverse direction to solve our task objective: by identifying additional qualified OOD latent encodings based on the established priors from a limited set of OOD images, we can synthesize unseen images without the need to adjust the original model parameters. Specifically, we note that the reconstruction ability of unseen data is subject to the prerequisite of the deterministic inversion (i.e., diffusion direction) and denoising (i.e., reserve direction) trajectories [48], as described in Sec. 2.3. To better understand the generalization abilities of DDPMs, we provide a theoretically grounded explanation to show that the inverted OOD encodings establish parameter-independent Gaussian priors, as detailed in Sec. 2.4. Meanwhile, we explicitly formulate the “mode separability” problem within the context of diffusion models, and empirically verify that the inverted OOD latent encodings are distinguishable from those corresponding to the training domain to avoid the “mode interference” issue in Sec. 2.5, which is also qualitatively illustrated in Fig. 3.
@@
Figure 2: Illustration of the trajectory-tuning based paradigm (left) and our proposed latent-discovery based paradigm (right) for OOD image synthesis with diffusion models. Given a pre-trained DM $p_{\theta}$ on images from domain $\mathcal{D}_{id}$, most existing methods seek to finetune the generation trajectories $p_{\theta^{\prime}}$ to synthesis data in a new domain $\mathcal{D}_{od}$. In contrast, we propose to discover unseen latent encodings to achieve the same goal via the frozen model $p_{\theta}$ by expanding the latent spaces.
The second contribution of this work introduces our paradigmatic design that achieves unseen domain synthesis in a tuning-free manner by discovering additional latent OOD encodings based on the inverted priors, as described in Sec. 3. Conceptually inspired by recent works that seek to manipulate the in-domain semantic attribute direction for data editing [74, 5], the key idea of our proposed OOD sampling method focuses on identifying potential latent directions by leveraging the geometric properties of the inverted OOD domains as additional domain-specific priors. The key takeaway from this part echoes our earlier analysis, demonstrating that in this latent discovery-based paradigm, the core technical challenge arises from the tendency of discovered OOD samples to be interfered with and captured by the original trained domain. This further underscores its distinguishable nature with intuitive tuning-based methods, where smaller data domain gaps are preferable for achieving better generalization performance.
As our last but not least contribution, we conduct extensive experiments in cross-model and cross-domain settings to demonstrate the effectiveness of our proposed paradigm, particularly its superiority in scenarios with large domain gaps between trained and target OOD data as detailed in Sec. 4. Notably, our experiments integrate both natural images (CelebA-HQ [26], AFHQ-Dog [8], LSUN-Church [65], LSUN-Bedroom [65]) and specific astrophysical data (GalaxyZoo [58] observation images and radiation field simulations of the dust emission [62]) to address a wide and representative range of domain gap differences. In addition to the quantitative and qualitative evaluation of natural images, widely adopted in the machine learning community, we also involve astrophysicists to assess the quality of generated data in comparison to their true astrophysical simulations. These comprehensive evaluations further reinforce our findings.
In summary, through the three contributions mentioned above, we aim to provide insights into a novel perspective for understanding and rethinking the domain generalization abilities of diffusion models, and to shed light on potential directions for broader interdisciplinary applications.
2 Domain Discovery and Expansion of DDPMs
This section presents the problem formulation of DDPMs [19] under the few-shot scenario and our analytical studies on the generalization properties of pre-trained DDPMs in the context of deterministic trajectories from both theoretical and empirical perspectives. The high-level takeaway from our in-depth study is that the inverted OOD samples establish Gaussian that is separable from the trained ID prior. The key technical challenge is to find qualified OOD latent free from the “mode interference”, which is distinguishable from the common understanding in tuning-based designs.
@@ 2.1 Problem Formulation
Given a DDPM $p_{\theta}$ trained on images from a domain $\mathcal{D}_{id}$, we aim to investigate the generalization properties of $p_{\theta}$ on other domain $\mathcal{D}_{od}$ using $N$ data samples $\mathbf{x}_{od}\in\mathcal{D}_{od}$, and eventually generate new data samples $\mathbf{x}_{od}^{\prime}\in\mathcal{D}_{od}$.
@@
The objective of DDPMs is similar to most previous generative models, which is to approximate an implicit data distribution $q(\mathbf{x}_{0})$ with a learned model distribution $p_{\theta}(\mathbf{x}_{0})$, as well as providing an easy-to-sample proxy (e.g., standard Gaussian). We further use $p_{s}$ and $p_{i}$ to represent the stochastic [19] and deterministic [48] generation processes, respectively. For the opposite direction, the pre-defined diffusion procedure is often denoted by $q(\mathbf{x}_{1:T}|q_{0})$. In addition, we use the hyper-parameter $\eta$ [48] to characterize the degree of stochasticity in the generative process, with $\eta=1$ for $p_{s}$ and $\eta=0$ for $p_{i}$. At intermediate stochastic levels, we adopt the notation $p_{\eta=k}$ with $k$ equals a constant between 0 and 1. Similar to existing literature, $T$ denotes the total diffusion steps. We use $\mathcal{X}_{t}$ to represent the latent (noisy) spaces formed by $\mathbf{x}_{t}$ along denoising.
2.2 Representation Ability of Latent Spaces
@@
We observe that a DDPM, trained even on a single-domain small dataset (e.g., dog faces), already has sufficient representation ability to accurately reconstruct arbitrary unseen images (e.g., human, church, and astrophysical data), as shown in the second column of Fig. 1. The reconstruction ability is subject to the deterministic inversion and denoising trajectories [48], which can be considered as a special case of vanilla stochastic process [19] as detailed in Sec. 2.3. The findings above suggest that: with a good mapping approximator (i.e., pre-trained DDPM) and proper tool (i.e., deterministic trajectories with DDIMs), its intermediate latent spaces already have sufficient representation ability for arbitrary images, which opens up the possibility to leverage DDPMs for synthesizing images from new domains without tuning the model parameters. The quantitative evaluation of the reconstruction results is in Tab. 3 in the Appendices.
2.3 Deterministic Diffusion Models
The core idea in the context of deterministic non-Markovian DDIMs [48] is to consider a family of $\mathcal{Q}$ of inference distributions, indexed by a real vector $\sigma\in\mathbb{R}^{T}_{\geq 0}$:
where $q_{\sigma}(\mathbf{x}_{T}|\mathbf{x}_{0})=\mathcal{N}(\sqrt{\alpha_{T}}\mathbf{x}_{0},(1-\alpha_{T})\mathbf{I})$ and $\forall t>1$,
$q_{\sigma}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})=\mathcal{N}(\sqrt{\alpha_{t-1}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t-1}-\sigma^{2}_{t}}\cdot\frac{\mathbf{x}_{t}-\sqrt{\alpha_{t}}\mathbf{x}_{0}}{\sqrt{1-\alpha_{t}}},\sigma^{2}_{t}\mathbf{I}),$ (2)
with $\alpha_{1:T}\in(0,1]^{T}$ is a decreasing sequence that parameterizes Gaussian transition kernels.
While this deterministic line of works was initially proposed to accelerate the vanilla ancestral sampling, later studies [31, 74, 64] revealed that the deterministic diffusion can be used as a tractable and lossless way for conducting data inversion, by connecting the above Eqn. 2 to neural ODEs. Similar deterministic variant is also proposed for the score-based formulations [51]. In this work, we leverage the deterministic properties as the tool for inversion and denoising the unseen images. More details about the deterministic inversion and denoising methods are included in Appendix C.
2.4 Parameter-Independent Properties: Gaussian Priors
We seek a theoretically grounded explanation to the generalization properties of pre-trained DDPMs after the inversion. The takeaway message is: In theory, the inverted latent encodings also establish Gaussian priors as presented in Lemma 3.444However, in practice, due to the fact that pre-trained DMs themselves are function approximators, the samples after inversion do not establish perfect Gaussians but rather approximations, echoing Sec. 3.
Lemma 2.1.
As also mentioned in [48], one can derive Lemma 3 by assuming for any $t\leq T$, $q_{\sigma}(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\sqrt{\alpha_{t}}\mathbf{x}_{0},(1-\alpha_{t})\mathbf{I})$ holds, if:
and then prove the statement with an induction argument for $t$ from $T$ to 1, since the base case $(t=T)$ already holds by definition. Proof details can be found in Appendix D, and we note that derivations are completed in the forward diffusion direction (i.e., the inversion direction from data to latent spaces), and make no modification to the trained model. This sets the primary rationale for estimating the latent distributions as Gaussian in a model parameter-independent manner.
2.5 Data-Dependent Properties: Mode Interference and Separability
In the literature of GANs-based generative models [15], “mode collapse” is a common issue that describes the training failure when generated images tend to be very similar given randomly sampled starting encodings from the Gaussian prior. Within the context of diffusion models in our work, we explicitly reveal a phenomenon analog to the “mode collapse” in GANs, which we refer to as “mode interference”, as qualitatively illustrated in Fig. 3 (a).
Notably, a key precondition to achieving the effective OOD latent sampling is that the established OOD prior mode should be separable from the ID Gaussian prior mode (i.e., a standard Gaussian). Otherwise, the denoised image would fall into the training domain as in Fig. 3 (b). The separability is further supported and validated by our empirical verification below in Sec. 2.6.
2.6 Analytical Experiments
We show empirical verification from multiple perspectives to support our parameter-independent and data-dependent properties described in Sec. 2.4 and Sec. 2.5.
Geometrical Properties of Gaussians. We leverage the geometrical measurements established of the high-dimensional studies in mathematics [7], as additional empirical support for the Gaussian priors in Sec. 2.4. Specifically, we compute several geometric metrics, including the pair-wise angles (angles formed by three arbitrary samples), sample-to-origin angles (angles formed by two arbitrary samples and the origin), pair-wise distance (euclidean distance between two arbitrary samples) and distance between OOD and ID Gaussian centers, and list the results in Tab. 1.
Characteristics above are typical geometric properties possessed by isotropic high-dimensional Gaussians [7]. Notably, three randomly sampled points from a high-dimensional Gaussian are almost surely form an equilateral triangle and are almost surely nearly orthogonal, which corresponds to the constant 60 pair-wise angle and 90 sample-to-origin angle in the first and second rows of Tab. 1, respectively. More details about the geometric properties are in Appendix D.
Mode Separability. As revealed by our analysis in Sec. 2.5, the separability between ID and OOD Gaussian modes is critical for synthesizing target unseen domain images without modifying the model parameters and for avoiding the “mode interference”. We further provide validation from the statistical and learning-based classifier perspectives to support the separability claim.
Statistical Validation. The separability of high-dimensional Gaussians follows Lemma 2.2 [7], which states that spherical Gaussians can be relaxingly separated by $\Omega(d^{\frac{1}{4}})$, or even $\Omega(1)$ with more sophisticated algorithms. In other words, for a DDPM trained on 256$\times$ 256 images with dimensionality $d=3\times 256\times 256$, ID and OOD modes can be well separated and avoid interference given a distance larger than $d^{\frac{1}{4}}\approx 21$, which is further validated by the empirical distance between centers, listed in the forth row of Tab. 1. More details about Lemma 2.2 are in Appendix D.
Lemma 2.2.
Mixtures of spherical Gaussians in $d$ dimensions can be separated provided their centers are separated by more than $d^{\frac{1}{4}}$ distance (i.e., a separation of $\Omega(d^{\frac{1}{4}})$).
Classifier Validation. Another empirical perspective to validate the separability between modes in the latent spaces is using the classifiers as in existing literature [45, 74]. Specifically, a linear classifier such as SVMs [17] can be fitted to test the separability between ID and OOD encodings in the latent spaces. In our analytical experiments, we fit SVMs on 1K inverted ID and OOD samples following the 7:3 training-testing ratio, and report the test accuracy in Tab. 1. As additional clarification, the classification results are obtained with the test on the latent space $\mathcal{X}_{T}$. Our rationale behind the choice of $T$ corresponds to the recent findings of DMs [74, 64], which indicates that $\mathcal{X}_{T}$, as the departure latent space, has the largest probabilistic support for the trained domain. In other words, if the latent ID and OOD modes can be separated in $\mathcal{X}_{T}$, they can be separated more easily in other $\mathcal{X}_{t}$, for $t=\{T-1,...,t,...1\}$.
Table 1: Geometric properties of inverted ID and OOD latent encodings. The results are computed based on 1K sample pairs. We report the mean and std for each geometric measurement to ensure the statistical significance. The base model is trained on AFHQ-Dog [8] in 256x256.
Pair-Angle 60.0$\pm$0 60.0$\pm$0 60.0$\pm$0.1 60.0$\pm$0.1 60.0$\pm$0.1 60.0 $\pm$ 0
Angle-Origin 89.7$\pm$0.01 89.7$\pm$0 89.8$\pm$0.01 89.7$\pm$0.01 89.1$\pm$0.01 87.6 $\pm$0.03
Pair-Distance 607.4$\pm$0.01 611.4$\pm$0.05 611.2$\pm$0.07 609.9$\pm$0.02 612.37$\pm$0.05 609.13 $\pm$ 0.1
Center-Distance - 33.3 26.0 33.2 54.7 60.8
Clf. Acc. - 0.97 0.99 0.99 1.0 1.0
As shown in Tab. 1, while performing the binary classification task on the inverted ID and OOD latent encodings, a simple linear classifier can well separate different domains, which further validates the latent modes are separable. In addition, we also observe that while the unseen images (e.g., human faces) are visually more similar to the trained domains (e.g., dogs), the inverted latent encodings inherit such similarity, making those unseen domains more difficult to be separated from the trained mode, and subsequently cause extra generation difficulties for those target domains via our tuning-free paradigm (see Sec. 3). We note this is distinguishable from previous tuning-based generalization works [69, 70, 56], which believe that it is usually easier to generalize model abilities to unseen domains similar to the trained ones, further validated in Sec. 4.
Our findings on the OOD mode separability also align with another thread of recent works that investigate pre-trained DDMs for discriminative tasks like classification and segmentation [32, 9, 38], where they reveal that diffusion models generalize better to classifying out-of-distribution images.
3 Discovery-based OOD Synthesis
Following our extensive analysis, we note that the methodological challenges in this work can be disentangled into two key points: sample qualified latent encodings from the OOD prior, and avoid the mode interference. Sampling from the inverted high-dimensional OOD priors, however, is an open and non-trivial challenging task given the theory-practice gap, [^5] As all the generative models can be considered as function approximators between the sampling prior and the implicit data distributions with certain error levels. and points to multiple possible directions of solutions such as classic statistical techniques. In fact, high-dimensional Gaussian estimation itself remains a challenging and complex research topic, especially in a multi-variant case [71, 4]. While we present our proposed latent sampling method below, we have experimented with many other methods that may not yield the best performance, such as vanilla Gaussian sampling and MCMC, with detailed discussions in Appendix E.
Few-Shot Latent References. Having obtained $\mathbf{x}_{od,t}$ from $N$ raw images $\mathbf{x}_{od}$, it seems intuitive to directly sample from the estimated Gaussian $\mathcal{N}(\mu_{est},\sigma^{2}_{est})$ from those inverted OOD latent encodings. However, we note this is insufficient to avoid the mode interference due to the fact that the empirical inversion does not establish perfect Gaussian in practice despite the theoretical Gaussian groundings, even for trained in-domain cases [74]. This is also qualitatively demonstrated in Fig. 3(c), where the inverted latent encodings preserve slight perceptible low-level visual features. To set a rather precise starting point in the OOD latent space, we thus propose to adopt the inverted samples $\mathbf{x}_{od,t}$ as the starting point to navigate the following sampling process.
Latent Direction for New Samples. While the inverted latent sample $\mathbf{x}_{od,t}$ serves as a known starting point, we continue to mine more unknown latent samples that shall lead to new denoised images $\mathbf{x}^{\prime}_{od}$. Specifically, inspired by several recent works in image editing through latent direction guidance [74, 5], we deploy the samples $\bar{\mathbf{x}}_{od,t}\sim\mathcal{N}(\mu_{est},\sigma^{2}_{est})$ drawn from the estimate Gaussian as the ultimate latent directions, and obtain samples $x^{\prime}_{od,t}$ along the spherical interpolation (slerp) [46] between $\mathbf{x}_{od,t}$ and $\bar{\mathbf{x}}_{od,t}$. The rationale behind the spherical interpolation comes from the fact that the probabilistic concentration mass of a high-dimensional Gaussian is mainly centered around a thin annulus around the equator [7]. In the meantime, it is critical for discovered latent samples to stay within the area of high probabilistic concentration mass to ensure denoised data with high quality.
Geometrical Optimization. So far we have localized a trajectory with intermediate samples $\mathbf{x}^{\prime}_{od,t}\in\small{\textsf{Slerp}}(\mathbf{x}_{od,t},\bar{\mathbf{x}}_{od,t})$ connecting two samples $\mathbf{x}_{od,t}$ and $\bar{\mathbf{x}}_{od,t}$ in this latent OOD space. To further improve the quality of our discovered latent samples, we leverage the geometric properties as domain-specific information to optimize the latent samples we have obtained from the previous step as additional criteria to avoid mode interference. Specifically, we can filter and reject a fraction of initial samples via the angles and distances as shown in Tab. 1 by setting pre-defined tolerance ranges $\omega$.
Overall Algorithm. The overall pipeline of our proposed OOD sampling method includes the following major steps: raw image inversion via $p_{i}$, geometric property computation (optional), latent sampling, and deterministic denoising via $p_{i}$. Details are depicted in Algo. 1 and 2 in Appendix E.
4 Generation Experiments

### v3

# Frontiers in Probabilistic Inference: Sampling Meets Learning

@@ Abstract
In this work, we investigate the domain generalization capabilities of diffusion models in the context of synthesizing images that are distinct from the training data. Instead of fine-tuning, we tackle this challenge from a sampling-based perspective using frozen, pre-trained diffusion models. Specifically, we demonstrate that arbitrary out-of-domain (OOD) images establish Gaussian priors in the latent spaces of a given model after inversion, and that these priors are separable from those of the original training domain. This OOD latent property allows us to synthesize new images of the target unseen domain by discovering qualified OOD latent encodings in the inverted noisy spaces, without altering the pre-trained models. Our cross-model and cross-domain experiments show that the proposed sampling-based method can expand the latent space and generate unseen images without impairing the generation quality of the original domain. We also showcase a practical application of our approach using astrophysical data, highlighting the potential of this generalization paradigm in data-sparse fields such as scientific exploration.
@@ 1 Introduction
Generalization ability, which enables a model to synthesize data from diverse domains, has long been a challenge for deep generative models. The current research trend focuses on leveraging larger models with more training data to facilitate improved generalization. The popularity of recent large-scale models, such as DALLE-2 [ramesh2022dalle2], Imagen [ho2022imagen], and StableDiffusion [stablediff], has demonstrated the impressive and promising representation capabilities of state-of-the-art (SOTA) diffusion generative models when trained on enormous image datasets. However, scaling up is not a panacea and does not fundamentally solve the generalization challenge. In other words, for data domains that remain sparse in these already giant datasets, such as astrophysical observation and simulation data, even SOTA models fail to synthesize data suitable for rigorous scientific research. In addition, scaling up requires extensive resources, severely limiting the number of research groups that are able to participate and contribute, and consequently hindering research progress. Given these concerns, our work focuses on studying generalization ability in a few-shot setup, where a pre-trained diffusion generative model and a small set of raw images different from its training domain are provided, with the ultimate objective of generating new data samples from the target OOD domain.
@@
Figure 1: Illustration of the trajectory-tuning based paradigm (left) and our proposed latent-sampling based paradigm (right) for OOD image synthesis with diffusion models. Given a pre-trained DM $p_{\theta}$ on images from domain $\mathcal{D}_{id}$, most existing methods seek to finetune the generation trajectories $p_{\theta^{\prime}}$ to synthesis data in a new domain $\mathcal{D}_{od}$. In contrast, we propose to discover unseen latent encodings to achieve the same goal via the frozen model $p_{\theta}$ by expanding the latent spaces.
To achieve this objective, one of the most straightforward and intuitive approaches is to fine-tune the pre-trained model using the target OOD images [kim2022diffusionclip, zhao2022egsde, kwon2022diffusion]. However, tuning-based methods have several drawbacks, particularly when there is a large gap between the training and target data domains. It is well established that tuning methods do not generalize well when the domain shift becomes too large [zhou2020learning, zhou2021domain, wang2019few]. In fact, our experiments show that fine-tuning pre-trained DMs using only image supervision with the vanilla variational lower bound loss [ho2020dpm] is very difficult; performance improves only when additional semantic guidance, such as the CLIP loss [radford2021clip], is introduced. Moreover, modifying model parameters can degrade synthesis quality on the original training domain, and the computational cost of tuning fully depends on the pre-trained model, which can be substantial given the well-known expense of training diffusion models.
In this work, we propose a sampling-based alternative to fine-tuning for tackling this problem. Our core insight lies in reframing the generalization challenge from “learning a new mapping function” to “discovering new OOD latents,” as illustrated in Fig.1. More specifically, we show that, after inversion, unseen OOD images exhibit several latent-space properties, including approximate Gaussianity and separability from the original training prior, as detailed in Sec.2. The former allows for the discovery of new OOD latent encodings using relatively simple sampling techniques, while the latter ensures that new OOD samples can be generated without interference from the original generation trajectories. We validate the effectiveness of this approach through a series of experiments.
2 Problem Formulation and Method
@@ 2.1 Problem Formulation
Given a diffusion denoising probablistic model (DDPM) $p_{\theta}$ trained on images from a domain $\mathcal{D}_{id}$, we aim to investigate the generalization properties of $p_{\theta}$ on other domain $\mathcal{D}_{od}$ using $N$ data samples $\mathbf{x}_{od}\in\mathcal{D}_{od}$, and eventually generate new data samples $\mathbf{x}_{od}^{\prime}\in\mathcal{D}_{od}$. The objective of DDPMs is similar to most previous generative models, which is to approximate an implicit data distribution $q(\mathbf{x}_{0})$ with a learned model distribution $p_{\theta}(\mathbf{x}_{0})$, as well as providing an easy-to-sample proxy (e.g., standard Gaussian). We further use $p_{s}$ and $p_{i}$ to represent the stochastic [ho2020dpm] and deterministic [song2020ddim] generation processes, respectively. For the opposite direction, the pre-defined diffusion procedure is often denoted by $q(\mathbf{x}_{1:T}|q_{0})$. Similar to existing literature, $T$ denotes the total diffusion steps. We use $\mathcal{X}_{t}$ to represent the latent (noisy) spaces formed by $\mathbf{x}_{t}$ along denoising.
2.2 Latent Sampling to New Domain Generalization
@@
Latent Representation of Unseen Images. We observe that a DDPM, trained even on a single-domain small dataset (e.g., dog faces), already has sufficient representation ability to accurately reconstruct arbitrary unseen images (e.g., human, church, and astrophysical data), as shown in Fig. 2. The reconstruction ability is subject to the deterministic inversion and denoising trajectories [song2020ddim]. The findings above suggest that: with a good mapping approximator (i.e., pre-trained DDPM) and proper tool (i.e., deterministic trajectories with DDIMs), its intermediate latent spaces already have sufficient representation ability for arbitrary images, which opens up the possibility to leverage DDPMs for synthesizing images from new domains without tuning the model parameters. Details about the deterministic inversion and reconstruction test can be found in the Appendix C.
Separability of Latent Gaussians. We then present two key properties of the latent OOD distributions after inversion, which are critical for developing an effective sampling-based method for domain generalization. First, the inverted OOD latent encoding exhibits (approximate) Gaussians in the latent spaces. The Gaussianity is established with both theoretical and empirical groundings, with details about these proofs included in the appendix. Second, the OOD priors are separable from the original Gaussian of the pre-trained image domains after the same inversion. We also show statistical support for this second property and its significance in the Appendix C.
Latent Sampling from OOD Priors. With the above properties, we then propose to generate new target images by first sampling from the OOD priors, and then denoise the newly sampled noisy latents via the deterministic DDIM denoising trajectories. Under the approximate Gaussian condition, the sampling is rather straightforward using the interpolation between arbitrary inverted OOD latents. Similarly, due to space constraints, we present additional details in the Appendix D.
Figure 2: Examples of synthesized OOD images across data domains and methods. (a) All the OOD samples are obtained via our proposed sampling method using a pre-trained DMs on AFHQ-Dog [choi2020stargan]. (b) The same sampling method can also be applied to achieve style transfer (e.g., between RGB and sketch images).
3 Experiments and Analysis